MONT 105Q – Mathematical Journeys Advice and Study Questions on Hamlet April 14, 2016

Advice

Because Shakespeare’s language is the English of his time (roughly 1580 - 1610 CE), and he was a master of creating picturesque effects with that language, the play contains many unusual and archaic words. In addition, meanings of familiar-looking words can be somewhat different from what you might expect. It can take some work even to decipher what Shakespeare means at some points. And when you do decipher the local meaning, it’s unfortunately very easy to overlook very important aspects of the plot of this play since they are often treated very briefly.

a. The Folger Shakespeare Library edition we are using has scene-by-scene synopses, excellent notes on unfamiliar words, complicated sentences, etc. printed on alternate pages facing the text of the play. Be sure you are following along with those as you read! From time to time, those notes will refer to a section of longer notes printed after the text of the play as well. Read those longer notes too.

b. Along with Prof. Isser in our CHQ cluster faculty (who’s directing the production of the play we’ll see next week), I highly recommend that you use the online summary and analysis of the play given at

           http://www.sparknotes.com/shakespeare/hamlet/

as you read, to supplement the notes in the Folger Shakespeare edition and to check your understanding.

Some Questions to Think About

0. Exactly who are Claudius, Gertrude, Polonius, Laertes, Horatio, Fortinbras, Ophelia, Rosenkrantz, Guildenstern? How are they related to Hamlet and to each other? Summarize what you can say about their characters in a few words or phrases.

1. Non-trivial, but perhaps “buried” plot points:

a. Hamlet refers to Wittenberg several times. What or who is he referring to? How

does this relate to Hamlet’s character and temperament? b. Exactly how many people does Hamlet kill or indirectly have killed by the end of

the play? Who are they? c. There’s an important journey that is a part of the plot but that is not directly

shown at all on stage. What is it? Who was on this journey? What was supposed to happen? What is the actual outcome?

d. How are the situations of Hamlet, Laertes, and Fortinbras all similar? e. Whose ghost appears several times in I? Why is he a ghost? What does he ask

Hamlet for in I.v? Does Hamlet understand why he is doing this? f. Shakespeare was probably the first playwright to use the device of showing a

play-within-a-play as part of the action. Where is that device used in Hamlet, what is the play, and what is the purpose of that in the action of this story? g. What’s the exact sequence of events in the climactic scene V.ii? h. Is Ophelia still alive at the end of the play? If not, exactly how does she die?

2. Suicide is a recurrent theme in Hamlet. How does the play treat the idea of suicide morally and religiously? Pay particular attention to the “O, that this too too solid flesh would melt” soliloquy (I.ii.129–158) and the “To be, or not to be” soliloquy (III.i.56– 88). Why does Hamlet believe that, although capable of suicide, most human beings choose to live despite the cruelty, pain, and injustice of the world? (And what does Horatio say about this that presents a somewhat different view?)

3. Does Hamlet have healthy relationships with the women in his world? First think about his relationship with his mother Gertrude during the time of the play (see for instance I.ii.144–163). Hamlet seems to be implying here and elsewhere that he thinks his mother was actually having an affair with Claudius, even before Claudius killed his father. Where does he get this idea? (look at I.v again!) Do you “read” Gertrude that way? Or is she just trying to preserve her status by latching onto Claudius once he has become King? Then think about Hamlets relationship with Ophelia. Does he love her during the main action of the play? Does he stop loving her at some point? Did he ever love her? Look at the “Get thee to a nunnery” scene, III.i, for instance. What do you make of that?

4. For a long time in the history of critical writing on Hamlet, it was common to take a deterministic view of Hamlet’s plot, arguing that the prince showed an inability to act and a tendency toward melancholy reflection that amounted to a “tragic flaw” leading inevitably to his demise. (Part of this is based on ideas about the Greek tragedies of Aeschylus, Euripides, and Sophocles, taken over into the analysis of Shakespeare’s tragedies.) Do you think that is an accurate way of understanding the play? Why or why not? Given Hamlets character and situation, would another outcome of the play have been possible?

5. Throughout the play, Hamlet claims to be feigning madness, but his portrayal of a madman is so intense and so convincing that many readers believe that Hamlet actually slips into insanity at certain moments in the play. Do you think this is true, or is Hamlet merely play-acting insanity? What evidence can you cite for either claim?

6. Look again at the gravediggers’ speeches in V.i, and at the speeches of Polonius and Osric. How are these episodes different from other sections of the play? What is Shakespeare doing there? What reaction are they supposed to produce? Do these scenes merely relieve the tension of the overtly tragic events, or do they serve a more serious thematic purpose as well? (Incidentally, a certain tradition of thought actually criticizes Shakespeare for including these scenes. According to ideas about tragedy inherited from Aristotle–tragedies were supposed to have unity of action, time, and place–some critics have argued that comedic elements should not appear in a tragedy. The classic 17th century French tragedies of Racine and Corneille, for instance, have nothing like Shakespeare’s use of comedic elements.)